# Exploiting Inter-Sample Correlation and Intra-Sample Redundancy for Partially Relevant Video Retrieval

### v1

@@ 3.3. Inter Correlation Enhancement
The PRVR task presents a fundamental semantic asymmetry, where textual queries only partially describe the content of target videos, leaving the visual modality inherently richer in semantic information. Our key observation reveals that while certain video moments may not align with their originally paired text annotations, they may exhibit strong semantic correspondence with other unpaired textual descriptions in the dataset as shown in Figure 1. This inter-sample correlation represents a valuable but underutilized resource in current PRVR approaches. We propose to exploit these underlying cross-modal relationships.
@@
Intuitionally, a truly correlated pair should not only mutually be the most similar to each other, but also exhibit high similarity. To further filter out mismatched noisy pairs, we pre-define $\beta_{t}$ as the correlation threshold. Only the pairs with a similarity higher than $\beta_{t}$ are selected.
@@ 3.4. Intra Redundancy Mining
The ICE module effectively harnesses inter-sample correlations to enhance the cross-modal semantic space for retrieval. However, while certain video semantics may align with unpaired text queries, they represent redundant information with respect to the originally paired text. This redundancy can introduce noise during video-text matching, thereby compromising retrieval accuracy.
To mitigate the intra-sample redundancy, we propose the Intra Redundancy Mining (IRM) module, which identifies redundant video moments by analyzing relationships among text, video, and moment-level features. These redundant features are then employed as hard negative samples during training. Unlike conventional negative samples (e.g., randomly sampled unpaired videos), these mined negatives exhibit higher feature similarity to the target moment, presenting a more challenging discrimination task. By forcing the model to distinguish these hard negatives, we promote the learning of more discriminative and query-relevant video representations, improving retrieval performance.
We construct redundant video features from two different perspectives. Technically, we model the semantic variation between the target moment feature $m^{t}$, video feature $v$, and query feature $q$ to obtain the video-view redundant feature $r_{v}\in\mathbb{R}^{D}$ and query-view redundant feature $r_{q}\in\mathbb{R}^{D}$:
(9) $r_{v}=FC(v-m^{t}),~{}r_{q}=FC(v-q),$
where $FC$ is a fully-connected layer to bridge the gap between video-moment and video-query differences. The intuition is that the video-level feature $v$ contains global semantics of all moments within the video. By subtracting the target moment feature $m^{t}$ from the video-level feature $v$, we obtain the video-view redundant feature $r_{v}$ which is irrelevant to the text. Similarly, we use a symmetrical operation to obtain the query-view redundant feature $r_{q}$ which is regarded to be irrelevant to the target moment. Since both redundant features $r_{v}$ and $r_{q}$ should be semantically irrelevant to the target moment feature and query feature, we employ $r_{v}$ and $r_{q}$ as negative samples in the moment-level similarity learning (Section 3.2.3), where the loss function is denoted as $\mathcal{L}_{neg}$.
Moreover, we further propose to directly align $r_{v}$ and $r_{q}$. As the two redundant features are obtained at the video-view and query-view respectively, their alignment boosts the learning of underlying relationships between videos and text queries in the cross-modal semantic space. The corresponding loss $\mathcal{L}_{red}$ is also the jointly application of Equation (5) and (6). The overall training objective of this module is the sum of the above two losses:
@@
Table 1. Comparison results on TVR, ActivityNet Captions, and Charades-STA. The best results are highlighted in bold, while the second-best outcomes are underlined. We achieve state-of-the-art results on all of the three datasets across all metrics.
@@
3.5. Temporal Coherence Prediction
Figure 4. The pipeline of the Temporal Coherence Prediction (TCP) module. The frame features are first divided into distinct groups, and group labels are assigned accordingly. Then, a subset of frame features is randomly selected and shuffled, and the group labels with these frames are predicted.
While our ICE and IRM modules effectively leverage inter-sample correlations and mitigate intra-sample redundancy, their performance fundamentally depends on the model’s ability to discern discriminative video features. Clearer feature distinctions enable more accurate cross-modal correlation measurement in ICE and better redundancy identification in IRM. To further enhance these capabilities, we introduce the Temporal Coherence Prediction (TCP) module, which strengthens temporal feature understanding through self-supervised learning. The pipeline is shown in Figure 4.
Specifically, the TCP module incorporates an auxiliary sequence prediction task. To capture the temporal structure of videos, we leverage the continuity between adjacent frames to partition the video sequence into coherent temporal segments. We group the video features and shuffle them to form an unordered sequence. The model is trained to predict the original temporal order. By explicitly modeling temporal coherence, this task helps the model to learn the coherence and causality between video events. Note that this strategy is applied to both the video-level branch and moment-level branch in Section 3.2.2 to enable multi-grained temporal coherence learning. We mainly take the video-level branch for explanation.
Concretely, we first group frame features into $g$ groups and each frame feature is given its group label ranging from $1$ to $g$. Every $T_{f}/g$ consecutive frame features are assigned with the same group label. Group labels of the original encoded frame features $V_{f}$ are denoted as $y_{o}\in\mathbb{R}^{T_{f}}$. We then randomly select and shuffle some frame features and encode them through the video encoder. Note that shuffling a small or large length of frames might lead to suboptimal performance. Too few shuffled frames provide insufficient learning signals, while excessive shuffling disrupts temporal coherence and makes the task overly challenging. Therefore, we pre-define $\beta_{r}$ as the shuffle ratio. The shuffled frame features and their shuffled group labels are represented as $\hat{V_{f}}=\left\{\hat{f_{i}}\right\}_{i=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times D}$ and $y_{s}\in\mathbb{R}^{T_{f}}$, respectively. Then we adopt a classifier to predict the group labels of both the original and shuffled frame features $V_{f}$ and $\hat{V_{f}}$:
(11) $p_{o}=Softmax(CLS(V_{f})),~{}p_{s}=Softmax(CLS(\hat{V_{f}})),$
where $CLS$ is a linear classifier, $p_{o}\in\mathbb{R}^{T_{f}\times g}$ and $p_{s}\in\mathbb{R}^{T_{f}\times g}$ are the predicted probability distributions of group labels. We adopt the cross-entropy loss to optimize this module:
(12) $\mathcal{L}_{tcp}=f_{CE}(p_{o},y_{o})+f_{CE}(p_{s},y_{s}),$
where $f_{CE}$ represents the cross-entropy loss function.
3.6. Training and Inference
Our proposed method is optimized by the loss of each module:
(13) $\mathcal{L}=\mathcal{L}_{base}+\lambda_{3}\mathcal{L}_{ice}+\lambda_{4}\mathcal{L}_{irm}+\lambda_{5}\mathcal{L}_{tcp},$
where $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are hyperparameters to balance losses.
During inference, the similarity of the given video-text pair $\mathcal{V}$ and $\mathcal{T}$ is the weighted sum of moment-level and video-level similarities following (Dong et al., 2022a; Wang et al., 2024b; Dong et al., 2023; Jiang et al., 2023; Yin et al., 2024):
@@ 4.1.1. Datasets
To evaluate the effectiveness of our method on the PRVR task, we utilize three untrimmed video datasets: TVR (Lei et al., 2020), ActivityNet Captions (Krishna et al., 2017), and Charades-STA (Gao et al., 2017). Note that we do not use moment-level annotations provided by these datasets in the setting of PRVR. Unlike conventional T2VR datasets (Xu et al., 2016; Chen and Dolan, 2011; Wang et al., 2019) using pre-trimmed short clips, PRVR employs untrimmed long videos where the content is only partially relevant to text queries.
@@ 4.1.3. Implementation Details
To ensure a fair comparison with previous works on PRVR (Dong et al., 2022a; Jiang et al., 2023; Dong et al., 2023; Wang et al., 2024b; Yin et al., 2024), we utilize the same visual and textual features provided by (Dong et al., 2022a). The visual and textual features are exacted by pre-trained I3D (Carreira and Zisserman, 2017) and RoBERTa (Liu et al., 2019), respectively. The feature dimension $D$ is set to 384. $T_{m}$, $m$, and $\alpha$ are set to 32, 0.2, and 0.7 following (Dong et al., 2022a; Wang et al., 2024b). $\beta_{t}$ and $\beta_{r}$ are pre-defined as 0.4 and 0.25. $g$ is set to 8. We set $\lambda_{\{1,2,3,4,5\}}$ to {0.02, 0.04, 0.1, 1, 1}. For model training, we utilize the Adam optimizer (Kingma and Ba, 2015) with an initial learning rate of 2.5e-4 and a mini-batch size of 128. The model is trained for 100 epochs with an early stopping strategy. Dropout (Srivastava et al., 2014) is set to 0.15 to prevent overfitting. All experiments are conducted on a single NVIDIA RTX 3090 GPU.
@@ 4.2. Comparison with State-of-the-Arts
We compare the proposed method with other state-of-the-art methods on TVR, ActivityNet Captions, and Charades-STA. These methods consist of both PRVR and T2VR models. The results are borrowed from (Dong et al., 2022a; Wang et al., 2024b) and the reported results of corresponding papers.
As summarized in Table 1, our method achieves state-of-the-art (SOTA) performance on all three datasets across all metrics. Specifically, we achieve significant improvements in SumR across these benchmarks: (1) On the TVR dataset, we outperform the previous SOTA method DL-DKD (Dong et al., 2023) by 14.1% absolute and 7.84% relative improvements. (2) On ActivityNet Captions, we surpass DL-DKD with gains of 12.8% absolute and 8.67% relative improvements. (3) On Charades-STA, we exceed PEAN (Jiang et al., 2023) by 6.6% absolute and 8.84% relative improvements. These results demonstrate the effectiveness and robustness of our approach.
4.3. Ablation Study
@@ 4.3.1. Main Ablation Studies
To comprehensively validate the effectiveness of each module in our method, we conduct extensive ablation studies in Table 2. In particular, Row 1 indicates the base model trained with $\mathcal{L}_{base}$ only, whose retrieval accuracy is far below the full model (Row 8) trained with Equation (13). Comparing the results in each row, we have the following observations:
(1) Each module independently improves performance over the base model (Row 1). The ICE module (Row 2) boosts SumR by 5.2, demonstrating its ability to enhance inter-sample semantic alignment. The IRM module (Row 3) improves SumR by 5.6, indicating its role in refining the cross-modal semantic space. The TCP module (Row 4) achieves the gain of 2.1 in SumR, highlighting its effectiveness in learning discriminative visual features.
@@
(3) Integrating all three modules leads to the best performance (Row 8), with significant gains across all metrics. This proves the effectiveness of jointly leveraging inter-sample correlations, intra-sample redundancy mining, and temporal coherence learning.
4.3.2. Inter Correlation Enhancement
We validate the efficacy of the correlation threshold $\beta_{t}$ in the Inter Correlation Enhancement module. As presented in Figure 5 (a), the choice of $\beta_{t}$ affects retrieval performance. When $\beta_{t}$ is set to 0.4, the model achieves the highest SumR of 81.3, outperforming both lower and higher thresholds. Lowering $\beta_{t}$ to 0.2 increases noise from weakly correlated pairs, reducing SumR to 80.1. On the other hand, raising $\beta_{t}$ to 0.6 or 0.8 over-suppresses potential correlations, thus degrading SumR to 79.6 and 78.6, respectively. Notably, the absence of $\beta_{t}$ (i.e., retaining all candidate pairs) yields a suboptimal SumR of 79.1, demonstrating that indiscriminate inclusion of low-similarity pairs could introduce harmful noise during training.
Figure 5. Ablation study of the correlation threshold $\beta_{t}$ in the Inter Correlation Enhancement module and the shuffle ratio $\beta_{r}$ in the Temporal Coherence Prediction module.
@@ 4.3.3. Intra Redundancy Mining
We investigate the effectiveness of each loss function and their combinations in the Intra Redundancy Mining (IRM) module. As shown in Table 3, both $\mathcal{L}_{neg}$ (Row 2) and $\mathcal{L}_{red}$ (Row 3) lead to improvements across all metrics. Moreover, combining $\mathcal{L}_{neg}$ and $\mathcal{L}_{red}$ yields the highest SumR, outperforming either loss alone. The effectiveness arises because $\mathcal{L}_{neg}$ directly suppresses redundancy by pushing away hard negatives, while $\mathcal{L}_{red}$ implicitly strengthens cross-modal alignment by correlating video-view and query-view redundancies.
4.3.4. Temporal Coherence Prediction
We assess the impact of the shuffle ratio $\beta_{r}$ in the Temporal Coherence Prediction (TCP) module. As illustrated in Figure 5 (b), the choice of $\beta_{r}$ influences the model’s ability to learn temporal coherence. When $\beta_{r}=0.25$, the model achieves the optimal SumR of 81.3, outperforming all other configurations. Adopting other ratios results in degradations in performance: smaller ratios yield insufficient learning signals, as minimal shuffling fails to challenge the model’s temporal reasoning. Conversely, larger ratios excessively disrupt temporal coherence, making the reordering task overly ambiguous and harming feature learning. Notably, the worst $\beta_{r}$ configuration even underperforms the full absence of TCP, demonstrating the necessity of setting a proper shuffle ratio.
We then demonstrate the contribution of applying the TCP module to both the video-level branch and the moment-level branch. As depicted in Table 4, leveraging TCP solely on the video-level branch (Row 2) improves SumR to 79.7, primarily enhancing fine-grained temporal modeling. Besides, applying TCP only to the moment-level branch (Row 3) achieves an improved SumR of 80.2, as moment-level coherence learning encourages the encoder to reason coarse-grained temporal relationships. Furthermore, the full setup (Row 4) integrates these advantages, achieving the best performance across all metrics. This dual-branch design ensures holistic temporal coherence learning at multiple granularities, which is essential for robust temporal feature learning.
4.3.5. Base Model
To demonstrate the broader generalization of our method, we further integrate it into two existing PRVR models: MS-SL (Dong et al., 2022a) and GMMFormer (Wang et al., 2024b). Concretely, we directly replace the base model with these models. As shown in Table 5, our approach acts as a plug-and-play enhancement to these models on the Charades-STA dataset. Consistent performance improvements are observed across both models. For instance, our method boosts SumR by 16.96% for MS-SL and 13.72% for GMMFormer in relative improvements, demonstrating its effectiveness and generalization. These results suggest that existing PRVR models inherently suffer from the limited cross-model alignment. In contrast, our method systematically exploits the inter-sample correlation and intra-sample redundancy, resulting in substantial performance improvements. The universal gains also highlight that our method is model-agnostic and effectively constructs a more comprehensive cross-modal semantic space.
@@ 4.3.5. Base Model
Figure 6. Visualization comparisons of retrieval results between our method, MS-SL (Dong et al., 2022a), and GMMFormer (Wang et al., 2024b) on ActivityNet Captions. Top-5 retrieval results are shown from left to right. Ground-truth videos are marked with red boxes. Zoom in for better visibility.
@@
Figure 7. t-SNE visualization (van der Maaten and Hinton, 2008) of text and video features on Charades-STA. (a) is the base model trained with only $\mathcal{L}_{base}$. (b) shows the full model trained with Equation (13).

### v2

@@ 3.3. Inter Correlation Enhancement
The PRVR task presents a fundamental semantic asymmetry, where textual queries only partially describe the content of target videos, leaving the visual modality inherently richer in semantic information. Our key observation reveals that while certain video moments may not align with their originally paired text annotations, they may exhibit strong semantic correspondence with other unpaired textual descriptions in the dataset, as demonstrated in Figure 1. This inter-sample correlation represents a valuable but unexplored resource in current PRVR approaches. We propose to exploit these underlying cross-modal relationships.
@@
Intuitively, a truly correlated pair should not only be the most similar to each other, but also exhibit high similarity. To further filter out mismatched noisy pairs and enhance the accuracy of correlated pairs, only the pairs with a similarity higher than the correlation threshold of 0.4 are selected.
@@ 3.4. Intra Redundancy Mining
The ICE module effectively harnesses inter-sample correlations to enhance the cross-modal semantic space for retrieval. However, while certain video semantics may align with unpaired text queries, they represent redundant information with respect to the originally paired text. This redundancy introduces noise during video-text matching, thereby limiting the retrieval accuracy.
To mitigate the intra-sample redundancy, we propose the Intra Redundancy Mining (IRM) module. It identifies redundant and query-irrelevant moment features by analyzing relations among text, video, and moment features. Unlike conventional negative samples (e.g., randomly sampled unpaired videos), these mined negatives exhibit higher feature similarity to the target moment, presenting a more challenging discrimination task. By forcing the model to distinguish redundant moments and query-relevant moments, we promote the learning of more discriminative and query-relevant video representations, improving retrieval performance.
We construct redundant moment features from two different perspectives. Technically, we model the semantic variation in the latent space (Wu et al., 2021; Qiu et al., 2025; Li et al., 2025) between the key moment feature $m^{k}$, video feature $v$, and query feature $q$ to obtain the video-view redundant feature $r_{v}\in\mathbb{R}^{D}$ and query-view redundant feature $r_{q}\in\mathbb{R}^{D}$:
(7) $r_{v}=FC(v-m^{k}),~r_{q}=FC(v-q),$
where $FC$ is a fully-connected layer to bridge the gap between video-moment and video-query differences. The intuition is that the video-level feature $v$ contains global semantics of all moments within the video. By subtracting the key moment feature $m^{k}$ from the video-level feature $v$, we obtain the video-view redundant feature $r_{v}$, which is deemed to be irrelevant to the text. Similarly, we use a symmetrical operation to obtain the query-view redundant feature $r_{q}$, which is regarded to be irrelevant to the target moment. Since both redundant features $r_{v}$ and $r_{q}$ should be semantically irrelevant to the key moment feature and query feature, we employ $r_{v}$ and $r_{q}$ as negative samples in the moment-level similarity learning (Section 3.2.3), where the loss function is denoted as $\mathcal{L}_{neg}$.
Moreover, we propose to directly align $r_{v}$ and $r_{q}$. As the two redundant features are obtained at the video-view and query-view respectively, their alignment boosts the learning of underlying relationships between videos and text queries in the cross-modal semantic space. The corresponding loss $\mathcal{L}_{red}$ is also the jointly application of Equation (3) and (4). The overall training objective of this module is the sum of the above two losses:
@@
3.5. Temporal Coherence Prediction
Figure 4. The pipeline of the Temporal Coherence Prediction (TCP) module. The frame features are first divided into distinct groups, and group labels are assigned accordingly. Then, a subset of frame features is randomly selected and shuffled, and the group labels of these frames are predicted.
While our ICE and IRM modules effectively leverage inter-sample correlations and mitigate intra-sample redundancy, their performance fundamentally depends on the model’s temporal modeling ability. Clearer temporal feature distinctions enable more accurate cross-modal correlation measurement in ICE and more precise redundancy identification in IRM. To further enhance these capabilities, we introduce the Temporal Coherence Prediction (TCP) module, which strengthens temporal structure learning through self-supervised learning. The pipeline is shown in Figure 4.
Specifically, the TCP module incorporates an auxiliary sequence prediction task. To capture the temporal structure of videos, we leverage the continuity between adjacent frames to divide the video sequence into coherent temporal segments. Inspired by HERO (Li et al., 2020), we shuffle the video features to form an unordered sequence. The model is trained to predict the original temporal order. By explicitly modeling temporal coherence, this task helps the model to learn the coherence and causality between video events. It is worth noting that, unlike HERO, which assigns extensive unique sequence labels, TCP adopts grouped labeling. We identify an issue in HERO’s labeling method. Predicting unique labels is overly hard and introduces noise. We address it by grouped labeling, reducing task complexity for robust temporal modeling. We also conduct ablations on group numbers in Section 4.3.3. Note that TCP is applied to both the video-level branch and moment-level branch in Section 3.2.2 to enable multi-grained temporal coherence learning. We mainly take the video-level branch for explanation.
Concretely, we first divide frame features into $g$ groups and each frame feature is given its group label ranging from $1$ to $g$. Every $T_{f}/g$ consecutive frame features are assigned the same group label. Group labels of the original encoded frame features $V_{f}$ are denoted as $y_{o}\in\mathbb{R}^{T_{f}}$. We then randomly select and shuffle some frame features and encode them through the video encoder. Note that shuffling a small or large length of frames might lead to suboptimal performance. Too few shuffled frames provide insufficient learning signals, while excessive shuffling disrupts temporal coherence and makes the task overly challenging. Therefore, we only shuffle 25% frames. The shuffled frame features and their shuffled group labels are represented as $\hat{V_{f}}=\left\{\hat{f_{i}}\right\}_{i=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times D}$ and $y_{s}\in\mathbb{R}^{T_{f}}$, respectively. Then we adopt a classifier to predict the group labels of both the original and shuffled frame features $V_{f}$ and $\hat{V_{f}}$:
(9) $),p_{o}=Softmax(CLS(V_{f})),~p_{s}=Softmax(CLS(\hat{V_{f}}$)),
where $CLS$ is the classifier, $p_{o}\in\mathbb{R}^{T_{f}\times g}$ and $p_{s}\in\mathbb{R}^{T_{f}\times g}$ are the predicted probability distributions of group labels. We adopt the cross-entropy loss to optimize this module:
(10) $\mathcal{L}_{tcp}=f_{CE}(p_{o},y_{o})+f_{CE}(p_{s},y_{s}),$
where $f_{CE}$ represents the cross-entropy loss function.
3.6. Inference
During inference, the ICE, IRM, and TCP modules are deactivated. The similarity of the given video-text pair $\mathcal{V}$ and $\mathcal{T}$ is the weighted sum of moment-level and video-level similarities in Section 3.2.3, following prior works (Dong et al., 2022a; Wang et al., 2024b; Dong et al., 2023; Jiang et al., 2023; Yin et al., 2024):
Table 1. Performance comparison results on TVR, ActivityNet Captions, and Charades-STA. The best results are highlighted in bold and the second-best outcomes are underlined. We achieve state-of-the-art results on all three datasets across all metrics. denotes using additional CLIP-B/32 model for knowledge distillation. indicates using Large Multimodal Models (LMMs).
@@ 4.1.1. Datasets
To evaluate the effectiveness of our method on the PRVR task, we utilize three widely-used datasets with untrimmed long videos: TVR (Lei et al., 2020), ActivityNet Captions (Krishna et al., 2017), and Charades-STA (Gao et al., 2017). Note that we do not use moment-level annotations provided by these datasets in the setting of PRVR. Unlike conventional T2VR datasets (Xu et al., 2016; Chen and Dolan, 2011; Wang et al., 2019) using pre-trimmed short clips, PRVR employs untrimmed long videos where the content is only partially relevant to text queries.
@@ 4.1.3. Implementation Details
For fair comparison with prior works on PRVR (Dong et al., 2022a; Jiang et al., 2023; Wang et al., 2024b; Yin et al., 2024; Dong et al., 2023; Zhang et al., 2025; Moon et al., 2025; Cho et al., 2025; Jun et al., 2025), we utilize the same visual and textual features provided by (Dong et al., 2022a). The visual and textual features are extracted by pre-trained I3D (Carreira and Zisserman, 2017) with ResNet152 (He et al., 2016) and RoBERTa (Liu et al., 2019). The feature dimension $D$ is set to 384. $\lambda_{1}$, $\lambda_{2}$, $T_{m}$, $m$, and $\alpha$ are set to 0.02, 0.04, 32, 0.2, and 0.7 following (Dong et al., 2022a; Wang et al., 2024b). $g$ is set to 8. For model training, we utilize the Adam optimizer (Kingma and Ba, 2015) with an initial learning rate of 2.5e-4 and a mini-batch size of 128. The model is trained for 100 epochs with an early stopping strategy. All experiments are conducted on a single NVIDIA RTX 3090 GPU. More implementation details are provided in the appendix.
@@ 4.2. Comparison with State-of-the-Arts
4.2.1. Compared Methods
We compare the proposed method with other state-of-the-art methods on TVR, ActivityNet Captions, and Charades-STA. These methods include PRVR (Dong et al., 2022a; Jiang et al., 2023; Wang et al., 2024b; Yin et al., 2024; Dong et al., 2023; Zhang et al., 2025; Moon et al., 2025; Cho et al., 2025; Jun et al., 2025), T2VR (Dong et al., 2021, 2022b; Wu et al., 2023; Li et al., 2023), and large multimodal models (Wang et al., 2024a). The results are sourced from (Dong et al., 2022a; Wang et al., 2024b) and corresponding papers.
4.2.2. Performance
As summarized in Table 1, our method achieves state-of-the-art (SOTA) results on all three datasets across all metrics. Notably, we outperform the previous SOTA method MGAKD (Zhang et al., 2025) without using additional CLIP models. Besides, we also surpass the Large Multimodal Model (LMM) InternVideo2 (Wang et al., 2024a) by a significant margin. InternVideo2 merely considers the similarity between text queries and entire videos, neglecting the essential fine-grained understanding of moment events. In contrast, our method fully leverages the rich semantics of distinct moment events and learns to distinguish them. Comparison highlights the advantages and expertise of our method in the PRVR task.
4.2.3. Efficiency
We evaluate the efficiency of our proposed model and existing models in the inference phase. For fair comparison, all experiments are conducted using the same batch size and a single RTX 3090 GPU on ActivityNet Captions. As shown in Table 2, with the same backbone and fewer learnable parameters, we achieve faster inference speed and better retrieval accuracy. These results demonstrate the effectiveness and efficiency of our proposed model.
Table 2. Efficiency comparison on ActivityNet Captions. All experiments are conducted under identical conditions.
Method Backbone Model Size Run-Time SumR
MS-SL (Dong et al., 2022a) I3D + RoBERTa 3.37M 146.63s 140.1
GMMFormer (Wang et al., 2024b) I3D + RoBERTa 11.37M 87.69s 146.0
Ours I3D + RoBERTa 4.65M 83.01s 160.4
4.3. Ablation Studies
@@ 4.3.1. Main Ablation Studies
To comprehensively validate the effectiveness of each module in our method, we conduct extensive ablation studies in Table 3. Comparing the results in each row, we have the following observations:
(1) Each module independently improves performance over the base model (Row 1-4), suggesting their effectiveness on refining the cross-modal semantic space.
@@
(3) Integrating all three modules leads to the best performance (Row 8), with significant gains across all metrics. In particular, TCP further improves ICE and IRM. ICE and IRM alone achieve a 9.1 absolute improvement in SumR (Row 1 and 7, 69.1 → 78.2). When combined with TCP, the performance gain increases to 10.1 (Row 4 and 8, 71.2 → 81.3).
@@ 4.3.2. Intra Redundancy Mining
We verify the effectiveness of each loss function in the Intra Redundancy Mining (IRM) module. As shown in Table 4, both $\mathcal{L}_{neg}$ (Row 2) and $\mathcal{L}_{red}$ (Row 3) lead to boosts on SumR. Moreover, combining $\mathcal{L}_{neg}$ and $\mathcal{L}_{red}$ (Row 4) yields the highest SumR. The effectiveness arises because $\mathcal{L}_{neg}$ directly suppresses redundancy by pushing away redundant and query-relevant moments, while $\mathcal{L}_{red}$ implicitly strengthens cross-modal alignment by correlating video-view and query-view redundancies.
Table 5. Ablations of the group number $g$ in the metric SumR.
Row $g$ Charades-STA TVR ActivityNet Captions
1 4 80.0 187.5 156.1
2 8 81.3 194.0 160.4
3 16 78.2 184.0 152.1
4 32 75.1 180.1 148.6
@@
Figure 5. Visualization comparisons of retrieval results between our method, MS-SL (Dong et al., 2022a), and GMMFormer (Wang et al., 2024b) on ActivityNet Captions. Top-5 retrieval results are shown from left to right. Ground-truth videos are marked with red boxes.
@@
Figure 6. t-SNE visualization (van der Maaten and Hinton, 2008) of text and video features on Charades-STA. (a) is the base model trained with only $\mathcal{L}_{base}$. (b) shows the full model trained with the full setup.
Figure 7. Text-moment similarity on TVR. The base model is trained with only $\mathcal{L}_{base}$, and the full model is trained with the full setup. We smooth out similarity intervals for better readability. Zoom in for better visibility.
4.3.3. Temporal Coherence Prediction
We assess the group number $g$ in the Temporal Coherence Prediction (TCP) module, which controls the granularity of temporal coherence learning. As shown in Table 5, performances decline as $g$ ¿ 8 in all datasets due to increased task difficulty. Notably, HERO’s method (Li et al., 2020) that assigns unique labels (i.e., $g$ = 32 for $T_{m}$ = 32 moment sequences) obtains the lowest results. The optimal results are achieved at $g$ = 8, suggesting that moderate grouping balances task difficulty and learning efficacy.
We then demonstrate the contribution of applying the TCP module to both the video-level branch and the moment-level branch. As depicted in Table 6, leveraging TCP solely on the video-level branch (Row 2) improves SumR to 79.7, primarily enhancing fine-grained temporal modeling. Besides, applying TCP only to the moment-level branch (Row 3) achieves an improved SumR of 80.2, as moment-level coherence learning encourages the encoder to reason coarse-grained temporal relationships. Furthermore, the full setup (Row 4) integrates these advantages, achieving the best performance across all metrics. This dual-branch design ensures holistic temporal coherence modeling at multiple granularities, which is essential for robust temporal structure learning.
4.3.4. Base Model
To demonstrate the generalization of our method, we further integrate it into two existing PRVR models: MS-SL (Dong et al., 2022a) and GMMFormer (Wang et al., 2024b). Concretely, we directly replace the base model with these models. As shown in Table 7, our approach acts as a plug-and-play enhancement to these models on the Charades-STA dataset. Consistent performance improvements are observed across both models. The universal gains also highlight that our method is model-agnostic and effectively constructs a more comprehensive cross-modal semantic space.